# Improved Margin Generalization Bounds for Voting Classifiers

### v1

@@ 2.1 New Margin Generalization Bounds
To remedy this, we exploit the clipping. We observe that for $f\in\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil}$, we have that $\sum_{x\in X}|f(x)|/m\leq\gamma_{1}$. We may thus invoke the result of [Rudelson and Vershynin, 2006] on the scaled function class $\gamma_{1}^{-1}\cdot\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil}=\{f\mid\exists f^{\prime}\in\operatorname{\Delta({\mathcal{H}})},f=\gamma_{1}^{-1}f^{\prime}\}$, i.e. the functions in $\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil}$ scaled by $\gamma_{1}^{-1}$ to get that
@@
$\ln{\left({\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil},\gamma_{0}/2)\right)}=O\left(d_{c\gamma_{0}\varepsilon/(2\gamma_{1})}\ln^{1+\varepsilon}{\left(\frac{m}{d_{c\gamma_{0}\varepsilon/(2\gamma_{1})}}\right)}\right),$
All that remains is thus to bound $d_{c\gamma_{0}\varepsilon/(2\gamma_{1})}$, the $\operatorname{\mathrm{fat}}_{c\gamma_{0}\varepsilon/(2\gamma_{1})}$-dimension of $\gamma_{1}^{-1}\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil}$ and then set $\varepsilon$ appropriate. Now the $\operatorname{\mathrm{fat}}_{c\gamma_{0}\varepsilon/(2\gamma_{1})}$-dimension of $\gamma_{1}^{-1}\cdot\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil}$ is, due to the scale invariance of $\operatorname{\mathrm{fat}}$-dimension, the same as the $\operatorname{\mathrm{fat}}_{c\gamma_{0}\varepsilon/2}$-dimension of $\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil}$. We have thus improved Eq. 13 to
@@ 2.2 Majority-of-3
We finally describe the main ideas in our proof that Majority-of-3 AdaBoosts achieves an optimal in-expectation error of $O(d/(\gamma^{2}m))$ as stated in Corollary 3. We will also explain why it is crucial for this result, that the logarithmic factors in our margin generalization bound Theorem 1 are $\ln(\gamma^{2}m/d)$ and not $\ln(m/d),\ln(m)$ or $\ln(\gamma m/d)$. Let ${\mathcal{D}}$ be the unknown data distribution over ${\mathcal{X}}$ and let $t\in\{-1,1\}^{\mathcal{X}}$ be the unknown target concept. Recall that if AdaBoost is run for $\Omega(\gamma^{-2}\ln m)$ iterations with a $\gamma$-weak learner ${\mathcal{W}}$, then it produces a voting classifier with margins $\Omega(\gamma)$ on all examples in the training sequence ${\mathbf{S}}\sim{\mathcal{D}}_{t}^{m}$. We now use an analysis idea by [Aden-Ali et al., 2024], used to show that the Majority-of-3 Empirical Risk Minimizers has an optimal in-expectation error for realizable PAC learning.
Consider partitioning a training sequence ${\mathbf{S}}\sim{\mathcal{D}}^{(2k-1)m}$ into $12k-$ 1 equal sized training sequences ${\mathbf{S}}_{1},\dots\mathbf{S}_{2k-1}$ (with $k=2$ for Majority-of-3 and $k=3$ for Majority-of-5) of $m$ training examples each (rescaling $m$ by $12k-$ 1 recovers the guarantees for a training sequence of size $m$). If we run AdaBoost on each to obtain voting classifiers $f_{{\mathbf{S}}_{1}},\dots,f_{{\mathbf{S}}_{2k-1}}$, then each $f_{{\mathbf{S}}_{i}}$ has margins $\Omega(\gamma)$ on all of ${\mathbf{S}}_{i}$. For concreteness, let us say all margins are at least $\gamma/2$. Furthermore, for any point $x\in{\mathcal{X}}$, we have that if the majority vote $\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},\cdots,f_{{\mathbf{S}}_{2k-1}})$ errs on $x$, where we define the majority vote as $\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},\ldots,f_{{\mathbf{S}}_{2k-1}})=\operatorname{\mathrm{sign}}(\sum_{i=1}^{2k-1}\operatorname{\mathrm{sign}}(f_{i}))$, then at least $k$ of the voting classifiers err on $x$. Let us fix an $x\in{\mathcal{X}}$ and denote by $p_{x}$ the probability that $f_{{\mathbf{S}}_{i}}$ errs on $x$, where the probability is over the random choice of ${\mathbf{S}}_{i}$. Observe that since the training sequences ${\mathbf{S}}_{i}$ are i.i.d., this is the same probability for each ${\mathbf{S}}_{i}$. Moreover, by independence of the ${\mathbf{S}}_{i}$’s, we have that the probability that a fixed set of $k$ of the voting classifiers all err on $x$ is precisely $p_{x}^{k}$. A union bound over all $\binom{2k-1}{k}\leq 2^{2k}$ choices of $k$ voting classifiers implies that
$\operatorname*{\mathbb{P}}_{{\mathbf{S}}}[\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},\cdots,f_{{\mathbf{S}}_{2k-1}})(x)\neq t(x)]\leq 2^{2k}p_{x}^{k}.$
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},\cdots,f_{{\mathbf{S}}_{2k-1}}))]$ $=\operatorname*{\mathbb{E}}_{({\mathbf{x}},t({\mathbf{x}}))\sim{\mathcal{D}}_{t}}[\operatorname*{\mathbb{P}}_{{\mathbf{S}}}[\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},\cdots,f_{{\mathbf{S}}_{2k-1}})\neq t({\mathbf{x}})]]$
@@
The goal is thus to bound $\operatorname*{\mathbb{P}}[R_{i}]$. For this, the key is to exploit that $p_{x}>2^{-i-1}$ for $x\in R_{i}$. Let $m_{i}=\operatorname*{\mathbb{P}}[R_{i}]m$ denote the expected number of samples from $R_{i}$ in a training sequence ${\mathbf{S}}_{j}\sim{\mathcal{D}}_{t}^{m}$. Since AdaBoost produces a voting classifier with margins $\Omega(\gamma)$ on all points in its training data set, it in particular has margins $\Omega(\gamma)$ on all data points in ${\mathbf{S}}_{j}\cap R_{i}$. We note that these samples are i.i.d. from the conditional distribution of a ${\mathbf{x}}\sim{\mathcal{D}}$, conditioned on ${\mathbf{x}}\in R_{i}$. Let us denote this conditional distribution by ${\mathcal{D}}_{t}\mid R_{i}$. We can now invoke our new margin generalization bound in Theorem 1 to conclude that
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{i}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}\mid R_{i}}(f_{{\mathbf{S}}_{i}})]=O\left(\frac{d\operatorname{\Gamma}(\gamma^{2}m_{i}/d)}{\gamma^{2}m_{i}}\right),$ (16)
with $\operatorname{\Gamma}(x)=\ln(x)\ln^{2}(\ln x)$. Note that Theorem 1 actually gives a high probability guarantee, which in particular implies the above guarantee on the expected error. The crucial point is that we invoke Theorem 1 with the conditional distribution ${\mathcal{D}}_{t}\mid R_{i}$ and $\operatorname{\mathcal{L}}_{{\mathbf{S}}_{i}}^{\gamma/2}(f_{{\mathbf{S}}_{i}})=0$ since we have all margins at least $\gamma/2$, and we have $m_{i}$ samples from this distribution (in expectation). On the other hand, we have by definition of $R_{i}$ that $\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{i}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}\mid R_{i}}(f_{{\mathbf{S}}_{i}})]>2^{-i-1}$. Writing $x=\gamma^{2}m_{i}/d$ for short, we thus conclude that
$\frac{\operatorname{\Gamma}(x)}{x}=\Omega\left(2^{-i}\right)\Rightarrow x=O\left(i\ln^{2}(i)2^{i}\right)\Rightarrow m_{i}=O\left(\frac{di\ln^{2}(i)2^{i}}{\gamma^{2}}\right)\Rightarrow\operatorname*{\mathbb{P}}[R_{i}]=O\left(\frac{di\ln^{2}(i)2^{i}}{\gamma^{2}m}\right)$
Inserting this in Eq. 15 bounds the expected error by (for constant $k$):
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},\cdots,f_{{\mathbf{S}}_{2k-1}}))]$ $=O\left(\sum_{i=0}^{\infty}\frac{2^{-ik}i\ln^{2}(i)2^{i}d}{\gamma^{2}m}\right).$ (17)
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{i}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}\mid R_{i}}(f_{{\mathbf{S}}_{i}})]=O\left(\frac{d\ln(\gamma m_{i}/d)}{\gamma^{2}m_{i}}\right),$ (18)
@@
Now plugging this bound $\operatorname*{\mathbb{P}}[R_{i}]$ into Eq. 15 (with $k=2$), we get the following error bounds for $\operatorname*{\mathbb{E}}_{{\mathbf{S}}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},f_{{\mathbf{S}}_{2}},f_{{\mathbf{S}}_{3}}))]$ of
@@
The obtained error bound of $\operatorname*{\mathbb{E}}_{{\mathbf{S}}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},f_{{\mathbf{S}}_{2}},f_{{\mathbf{S}}_{3}}))]$ thus has a superfluous $\ln{\left(1/\gamma\right)}$ factor.
These shortcomings of previous margins bounds used in the above analysis of the expected error $\operatorname*{\mathbb{E}}_{{\mathbf{S}}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},f_{{\mathbf{S}}_{2}},f_{{\mathbf{S}}_{3}}))]$, adding a superfluous $\ln{\left(1/\gamma\right)}$-factor, is precisely the reason why previous work needed a Majority-of-5. Being unable to use the margin generalization bounds with sub-optimal $\ln(\cdot)$ factors, the work [Høgsgaard et al., 2024] instead relied on the much weaker guarantee that any voting classifier $f$ with margins $\gamma$ has $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{t}}(f)=O(\sqrt{d/(\gamma^{2}m)}),$ where this bound can be obtained by following the steps of [Schapire and Freund, 2012] [page 107-111] and using the stronger bound on the Rademacher complexity for a function class with VC-dimension $d$ of $O(\sqrt{d/m})$ due to [Dudley, 1978] [See e.g. Theorem 7.2 [Hajek and Raginsky, 2021]], instead of the weaker $O(\sqrt{d\ln{\left(m/d\right)}/m})$ used in [Schapire and Freund, 2012]. This bound has the right behaviour for $m\approx d/\gamma^{2}$ unlike the bounds with sub-optimal logarithmic factors and results in the guarantee $\operatorname*{\mathbb{P}}[R_{i}]=O(2^{2i})$ instead of $O(i\ln^{2}(i)2^{i})$. This needs $k=3$ for $2^{-ik}$ to dominate $2^{2i}$ in $\sum_{i=0}^{\infty}2^{-ik}2^{2i}$ from  Eq. 17, whereas it suffices for us with $k=2$ since we only need to bound $\sum_{i=0}^{\infty}2^{-ik}i\ln^{2}(i)2^{i}$, leading to the more natural Majority-of-3 instead of Majority-of-5.

### v2

@@ 2.1 New Margin Generalization Bounds
To remedy this, we exploit the clipping. We observe that for $f\in\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil}$, we have that $\sum_{x\in X}|f(x)|/m\leq\gamma_{1}$. We may thus invoke the result of [Rudelson and Vershynin, 2006] on the scaled function class $\gamma_{1}^{-1}\cdot\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil}=\{f\mid\exists f^{\prime}\in\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil},f=\gamma_{1}^{-1}f^{\prime}\}$, i.e. the functions in $\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil}$ scaled by $\gamma_{1}^{-1}$ to get that
@@
$\ln{\left({\mathcal{N}}_{\infty}(X,\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil},\gamma_{0}/2)\right)}=O\left(d_{c\gamma_{0}\varepsilon/(2\gamma_{1})}\ln^{1+\varepsilon}{\left(\frac{m}{d_{c\gamma_{0}\varepsilon/(2\gamma_{1})}}\right)}\right).$
Now the $\operatorname{\mathrm{fat}}_{c\gamma_{0}\varepsilon/(2\gamma_{1})}$-dimension of $\gamma_{1}^{-1}\cdot\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil}$ is, due to the scale invariance of $\operatorname{\mathrm{fat}}$-dimension, the same as the $\operatorname{\mathrm{fat}}_{c\gamma_{0}\varepsilon/2}$-dimension of $\operatorname{\Delta({\mathcal{H}})}_{\left\lceil\gamma_{1}\right\rceil}$. We have thus improved Eq. 13 to
@@ 2.2 Majority-of-3
We finally describe the main ideas in our proof that Majority-of-3 AdaBoosts achieves an optimal in-expectation error of $O(d/(\gamma^{2}m))$ as stated in Corollary 3. We will also explain why it is crucial for this result, that the logarithmic factors in our margin generalization bound Theorem 1 are $\ln(\gamma^{2}m/d)$ and not $\ln(m/d),\ln(m)$ or $\ln(\gamma m/d)$. We remark that the later bounds can not be turned into $\ln(\gamma^{2}m/d)$ by multiplying by some constant factor, for instance $2\ln{\left(\gamma m/d\right)}=\ln{\left(\gamma^{2}m^{2}/d^{2}\right)}\not\leq\ln{\left(\gamma^{2}m/d\right)}$ Let ${\mathcal{D}}$ be the unknown data distribution over ${\mathcal{X}}$ and let $c^{*}\in\{-1,1\}^{\mathcal{X}}$ be the unknown target concept. Recall that if AdaBoost is run for $\Omega(\gamma^{-2}\ln m)$ iterations with a $\gamma$-weak learner ${\mathcal{W}}$, then it produces a voting classifier with margins $\Omega(\gamma)$ on all examples in the training sequence ${\mathbf{S}}\sim{\mathcal{D}}_{c^{*}}^{m}$. We now use an analysis idea by [Aden-Ali et al., 2024], used to show that the Majority-of-3 Empirical Risk Minimizers has an optimal in-expectation error for realizable PAC learning.
Consider partitioning a training sequence ${\mathbf{S}}\sim{\mathcal{D}}^{(2k-1)m}_{c^{*}}$ into $12k-$ 1 equal sized training sequences ${\mathbf{S}}_{1},\dots\mathbf{S}_{2k-1}$ (with $k=2$ for Majority-of-3 and $k=3$ for Majority-of-5) of $m$ training examples each (rescaling $m$ by $12k-$ 1 recovers the guarantees for a training sequence of size $m$). If we run AdaBoost on each to obtain voting classifiers $f_{{\mathbf{S}}_{1}},\dots,f_{{\mathbf{S}}_{2k-1}}$, then each $f_{{\mathbf{S}}_{i}}$ has margins $\Omega(\gamma)$ on all of ${\mathbf{S}}_{i}$. For concreteness, let us say all margins are at least $\gamma/2$. Furthermore, for any point $x\in{\mathcal{X}}$, we have that if the majority vote $\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},\cdots,f_{{\mathbf{S}}_{2k-1}})$ errs on $x$, where we define the majority vote as $\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},\ldots,f_{{\mathbf{S}}_{2k-1}})=\operatorname{\mathrm{sign}}(\sum_{i=1}^{2k-1}\operatorname{\mathrm{sign}}(f_{i}))$, then at least $k$ of the voting classifiers err on $x$. Let us fix an $x\in{\mathcal{X}}$ and denote by $p_{x}:=\operatorname*{\mathbb{P}}_{{\mathbf{S}}_{i}}[f_{{\mathbf{S}}_{i}}(x)\not=c^{*}(x)]$ the probability that $f_{{\mathbf{S}}_{i}}$ errs on $x$, where the probability is over the random choice of ${\mathbf{S}}_{i}$. Observe that since the training sequences ${\mathbf{S}}_{i}$ are i.i.d., this is the same probability for each ${\mathbf{S}}_{i}$. Moreover, by independence of the ${\mathbf{S}}_{i}$’s, we have that the probability that a fixed set of $k$ of the voting classifiers all err on $x$ is precisely $p_{x}^{k}$. A union bound over all $\binom{2k-1}{k}\leq 2^{2k}$ choices of $k$ voting classifiers implies that
$\operatorname*{\mathbb{P}}_{{\mathbf{S}}}[\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},\cdots,f_{{\mathbf{S}}_{2k-1}})(x)\neq c^{*}(x)]\leq 2^{2k}p_{x}^{k}.$
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},\cdots,f_{{\mathbf{S}}_{2k-1}}))]$ $=\operatorname*{\mathbb{E}}_{({\mathbf{x}},c^{*}({\mathbf{x}}))\sim{\mathcal{D}}_{c^{*}}}[\operatorname*{\mathbb{P}}_{{\mathbf{S}}}[\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},\cdots,f_{{\mathbf{S}}_{2k-1}})\neq c^{*}({\mathbf{x}})]]$
@@
The goal is thus to bound $\operatorname*{\mathbb{P}}[R_{i}]$. For this, the key is to exploit that $p_{x}>2^{-i-1}$ for $x\in R_{i}$. Let $m_{i}=\operatorname*{\mathbb{P}}[R_{i}]m$ denote the expected number of samples from $R_{i}$ in a training sequence ${\mathbf{S}}_{j}\sim{\mathcal{D}}_{c^{*}}^{m}$. Since AdaBoost produces a voting classifier with margins $\Omega(\gamma)$ on all points in its training data set, it in particular has margins $\Omega(\gamma)$ on all data points in ${\mathbf{S}}_{j}\cap R_{i}$. We note that these samples are i.i.d. from the conditional distribution of a ${\mathbf{x}}\sim{\mathcal{D}}$, conditioned on ${\mathbf{x}}\in R_{i}$. Let us denote this conditional distribution by ${\mathcal{D}}_{c^{*}}\mid R_{i}$. We can now invoke our new margin generalization bound in Theorem 1, and conclude that
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{j}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}\mid R_{i}}(f_{{\mathbf{S}}_{j}})]=O\left(\frac{d\operatorname{\Gamma}(\gamma^{2}m_{i}/d)}{\gamma^{2}m_{i}}\right),$ (16)
with $\operatorname{\Gamma}(x)=\ln(x)\ln^{2}(\ln x)$. Note that Theorem 1 actually gives a high probability guarantee, which by some calculations implies the above guarantee on the expected error. The crucial point in these calculations is that we invoke Theorem 1 with the conditional distribution ${\mathcal{D}}_{c^{*}}\mid R_{i}$ and $\operatorname{\mathcal{L}}_{{\mathbf{S}}_{j}\cap R_{i}}^{\gamma/2}(f_{{\mathbf{S}}_{j}})=0$ since we have all margins at least $\gamma/2$, and we have $m_{i}$ samples from this distribution (in expectation). On the other hand, we also have by definition of $R_{i}$, that Eq. 16 can be lower bouned as $\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{j}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}\mid R_{i}}(f_{{\mathbf{S}}_{j}})]>2^{-i-1}$. Writing $x=\gamma^{2}m_{i}/d$ for short, we thus conclude that
$\frac{\operatorname{\Gamma}(x)}{x}=\Omega\left(2^{-i}\right)\Rightarrow x=O\left(i\ln^{2}(i)2^{i}\right)\Rightarrow m_{i}=O\left(\frac{di\ln^{2}(i)2^{i}}{\gamma^{2}}\right)\Rightarrow\operatorname*{\mathbb{P}}[R_{i}]=O\left(\frac{di\ln^{2}(i)2^{i}}{\gamma^{2}m}\right),$
where the last equality uses that $m_{i}=\operatorname*{\mathbb{P}}[R_{i}]n$. Inserting this in Eq. 15 bounds the expected error by (for constant $k$):
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},\cdots,f_{{\mathbf{S}}_{2k-1}}))]$ $=O\left(\sum_{i=0}^{\infty}\frac{2^{-ik}i\ln^{2}(i)2^{i}d}{\gamma^{2}m}\right).$ (17)
@@
$\operatorname*{\mathbb{E}}_{{\mathbf{S}}_{i}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}\mid R_{i}}(f_{{\mathbf{S}}_{i}})]=O\left(\frac{d\ln(\gamma m_{i}/d)}{\gamma^{2}m_{i}}\right),$ (18)
@@
Now plugging this bound $\operatorname*{\mathbb{P}}[R_{i}]$ into Eq. 15 (with $k=2$), we get the following error bounds for $\operatorname*{\mathbb{E}}_{{\mathbf{S}}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},f_{{\mathbf{S}}_{2}},f_{{\mathbf{S}}_{3}}))]$ of
@@
The obtained error bound of $\operatorname*{\mathbb{E}}_{{\mathbf{S}}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},f_{{\mathbf{S}}_{2}},f_{{\mathbf{S}}_{3}}))]$ thus has a superfluous $\ln{\left(1/\gamma\right)}$ factor - if the bound in Eq. 18 where with $\ln{\left(\gamma m/d\right)}$ and $\ln{\left(m/d\right)}$ instead the above analysis would again give a superfluous factor.
These shortcomings of previous margins bounds used in the above analysis of the expected error $\operatorname*{\mathbb{E}}_{{\mathbf{S}}}[\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(\operatorname{\mathrm{Maj}}(f_{{\mathbf{S}}_{1}},f_{{\mathbf{S}}_{2}},f_{{\mathbf{S}}_{3}}))]$, introducing superfluous factors, is precisely the reason why previous work needed a Majority-of-5. Being unable to use the margin generalization bounds with sub-optimal $\ln(\cdot)$ factors, the work [Høgsgaard et al., 2024] instead relied on the much weaker guarantee that any voting classifier $f$ with margins $\gamma$ has $\operatorname{\mathcal{L}}_{{\mathcal{D}}_{c^{*}}}(f)=O(\sqrt{d/(\gamma^{2}m)}),$ where this bound can be obtained by following the steps of [Schapire and Freund, 2012] [page 107-111] and using the stronger bound on the Rademacher complexity for a function class with VC-dimension $d$ of $O(\sqrt{d/m})$ due to [Dudley, 1978] [See e.g. Theorem 7.2 [Hajek and Raginsky, 2021]], instead of the weaker $O(\sqrt{d\ln{\left(m/d\right)}/m})$ used in [Schapire and Freund, 2012]. This bound has the right behaviour for $m\approx d/\gamma^{2}$ unlike the bounds with sub-optimal logarithmic factors and results in the guarantee $\operatorname*{\mathbb{P}}_{{\mathbf{x}}\sim{\mathcal{D}}}[R_{i}]=O(d2^{2i}/(\gamma^{2}m))$ instead of $O(di\ln^{2}(i)2^{i}/(\gamma^{2}m))$. This needs $k=3$ for $2^{-ik}$ to dominate $2^{2i}$ in $\sum_{i=0}^{\infty}2^{-ik}2^{2i}$ from  Eq. 17, whereas it suffices for us with $k=2$ since we only need to bound $\sum_{i=0}^{\infty}2^{-ik}i\ln^{2}(i)2^{i}$, leading to the more natural Majority-of-3 instead of Majority-of-5.